OpenStudy (anonymous):

PLEASE HELP! FAN + MEDAL Which calculation can be used to find the value of p in the equation p^3 = 8?

Nnesha (nnesha):

what would you do to solve x^2 ?

OpenStudy (anonymous):

These are the options

OpenStudy (anonymous):

It's review and I forgot how to do this! If I get a bad grade in review my teach won't let me take the test :(

Nnesha (nnesha):

exponent rule is you can change root to exponent \[\huge\rm \sqrt[n]{x^m} = x^\frac{ m }{ n }\] so if you have to solve x^2 you would take square root and then apply exponent rule(change root to exponent )\[\large\rm x^2 = \sqrt{x^2} = x^\frac{ 2}{ 2 }\] both 2 can cancel out \[\large\rm x^\frac{ \cancel 2 }{ \cancel 2 }=x^1 =x\]

OpenStudy (anonymous):

I'm so confused

Nnesha (nnesha):

\[\large\rm x^4 = \sqrt[4]{x^4}\] \[\large\rm x^5 = \sqrt[5]{x^5}\] to cancel out exponent you should take root both side

OpenStudy (anonymous):

Ohhh

Nnesha (nnesha):

yes....